OpenStudy (anonymous):

Put the equation in standard form for a hyperbola. 9x^2-18x-25y^2-150Y=441

OpenStudy (anonymous):

9x^2-18x-25y^2-150y=441 =>factorising the x term and y term we get , 9(x-1)^2-25(y+3)^2=125

OpenStudy (anonymous):

check the calculation, please. eventually u need to bring the equation in for (x-a)^2/A-(y-b)^2 /B=1

OpenStudy (anonymous):

the factorisation i've done it for you.

OpenStudy (anonymous):

thats what I got _ yay!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

one sec i redid it and got a diff

OpenStudy (anonymous):

yea tell...

OpenStudy (anonymous):

- got 225 on the right side

OpenStudy (anonymous):

i think you just made a typo

OpenStudy (anonymous):

it's all good meow